24. While observing the obelisk in Saint Peter’s Square in Vatican City, you stand at the edge of the shadow it is casting across the ground, which is about 120 m long. While kneeling, you estimate your line of sight with the top of the obelisk to be about 12o above the horizontal. Approximately how tall is the obelisk? (3)

A) 9.6 m C) 23.1 m E) 26 m 

B) 120 m D) 190 m 



1
Expert's answer
2021-06-12T10:42:23-0400

24) We can find the height of the obelisk from the geometry:


"tan\theta=\dfrac{h}{L},""h=Ltan\theta=120\ m\cdot tan12^{\circ}=26\ m."

Therefore, the correct answer is (E) 26 m.